HIGH RESOLUTION THZ AND FIR SPECTROSCOPY OF SOCl_{2}

\underline{M. A. MARTIN-DRUMEL}, A. CUISSET, D. A. SADOVSKII, G. MOURET, F. HINDLE, Laboratoire de Physico-Chimie de l’Atmosphère, EA 4493, Université du Littoral Côte d’Opale, 59140 Dunkerque, France; O. PIRALI, Institut des Sciences Moléculaires d’Orsay, CNRS, UMR 8214, Université Paris XI, bât. 210, 91405 Orsay Cedex, France; SOLEIL Synchrotron, AILES beamline, L’orme des Merisiers, Saint-Aubin, 91192 Gif-Sur-Yvette, France.

Thionyl chloride (SOCl_{2}) is an extremely powerful oxidant widely used in industrial processes and playing a role in the chemistry of the atmosphere^{a}. In addition, it has a molecular configuration similar to that of phosgene (COCl_{2}), and is therefore of particular interest for security and defense applications. Low resolution vibrational spectra of gas phase SOCl_{2}^{b} as well as high resolution pure rotational transitions up to 25 GHz^{c} have previously been investigated. To date no high resolution data are reported at frequencies higher than 25 GHz.

We have investigated the THz absorption spectrum of SOCl_{2} in the spectral region 70–650 GHz using a frequency multiplier chain coupled to a 1 m long single path cell containing a pressure of about 15 μbar. At the time of the writing, about 8000 pure rotational transitions of SO^{35}Cl_{2} with highest J and K_{a} values of 110 and 50 respectively have been assigned on the spectrum. We have also recorded the high resolution FIR spectra of SOCl_{2} in the spectral range 50–700 cm^{−1} using synchrotron radiation at the AILES beamline of SOLEIL facility. A White-type cell aligned with an absorption path length of 150 m has been used to record, at a resolution of 0.001 cm^{−1}, two spectra at pressures of 5 and 56 μbar of SOCl_{2}. On these spectra all FIR modes of SOCl_{2} are observed (ν_{2} to ν_{6}) and present a resolved rotational structure. Their analysis is in progress.

aT. J. Johnson et al., J. Phys. Chem. A 107, 6183 (2003)

bD. E. Martz and R. T. Lagemann, J. Chem. Phys. 22,1193 (1954)

cH. S. P. Müller and M. C. L. Gerry, J. Chem. Soc. Faraday Trans. 90, 3473 (1994)